OpenStudy (anonymous):

Find 20% of 150. How would you do this problem..?

OpenStudy (anonymous):

150 x (20/100)

OpenStudy (anonymous):

= 150 x 1/5

Parth (parthkohli):

20% is the same as 0.20 or 0.2 'of' means multiply. \( \color{Black}{\Rightarrow 0.2 \times 150 }\)

OpenStudy (anonymous):

you do it in your head 150 ten percent is 15 twenty percent is 30

Parth (parthkohli):

People go so well but then they give the answer =.=

OpenStudy (anonymous):

I don't want the answer~

jhonyy9 (jhonyy9):

so you can make the calcule using the rule of three simple than you write 150 ---- 100% x ----- 20% -------------- so from this you calcule the x what will be equal x=(150*20)/100 =3000/100 =30 so x=30 - hope so much that you will can understanding this calcules good luck bye